### v1

@@ Abstract
Context. Merging compact objects such as binary black holes provide a promising probe for the physics of dark matter (DM). The gravitational waves emitted during inspiral potentially allow to detect DM spikes around black holes. This is because the dynamical friction force experienced by the inspiraling black hole alters the orbital period and thus the gravitational wave signal.
@@
Methods. To study the dynamical friction force, we use idealized $N$-body simulations, where we include self-interacting dark matter.
Results. We find that the dynamical friction force for inspiraling black holes would be typically enhanced by DM self-interactions compared to a collisionless medium (ignoring differences in the DM density). At lower velocities below the sound speed, we find that the dynamical friction force can be reduced by the presence of self-interactions.
@@ 1 Introduction
Since the first direct detection of gravitational waves (GWs) from merging stellar-mass black holes (BHs) and neutron stars with the LIGO detectors (Abbott et al., 2016, 2017a, 2017b, 2019), 100 years after the formulation of Einstein’s theory of General Relativity (Einstein, 1916) and about 40 years of the first indirect evidence (Hulse & Taylor, 1975; Taylor et al., 1976; Weisberg et al., 1981), GW astronomy has become an important and very promising field to probe fundamental physics. The GWs emitted by binary mergers do not only allow for testing general relativity at unprecedented levels, but also to probe other physical processes. This in particular includes the physics of neutron stars. Binary mergers of those objects could allow probing the equation of state of ultrahigh-density matter (e.g. Baiotti, 2022). In the future, GW detectors will become sensitive to sub-Hertz GW frequencies, thanks to space-based GW observatories such as LISA (Amaro-Seoane et al., 2017) and Taiji (Hu & Wu, 2017). This will allow us to observe the evolution of intermediate- and extreme-mass ratio inspirals (I/EMRIs), in which a small stellar-mass compact object inspirals into a much larger intermediate/supermassive black hole, over long times and provide a sensitive probe of BH environments.
@@
BHs might be “dressed” with DM, i.e., surrounded by a dense DM spike (e.g. Gondolo & Silk, 1999; Bertone et al., 2005; Bertone, 2024). The first robust evidence for an observation of a DM spike around a BH is claimed to be found by Chan & Lee (2024), see also the work by Alachkar et al. (2023). They studied the supermassive black hole binary system OJ 287 with outburst observations covering a time span of 104 years. The presence of a DM spike can impact the orbit of an inspiralling BH and make the GW signal sensitive to DM physics. This could in principle be detectable, e.g., for M87* (Daghigh & Kunstatter, 2024). The inspiralling BH is decelerated by the dynamical friction force, which depends on the density and the velocity dispersion of the DM spike (Chandrasekhar, 1943). In consequence, the orbital decay of an inspiralling BH allows probing properties of the DM spike and thus the particle nature of DM.
@@
Potentially, such DM spikes could lead to a faster decay of the orbit at small separations ($\approx 1$ pc). This would provide an explanation for the final parsec problem, i.e., how binary BHs can lose enough energy and angular momentum to reach separations small enough that the energy loss due to the emission of GWs is sufficient to make them merge within a sufficiently small timescale (e.g. Milosavljević & Merritt, 2003). In this context, Alonso-Álvarez et al. (2024) claim that collisionless DM is not sufficient to make the BH orbit decay fast enough because the DM spike would be destroyed by the friction energy injected from the BHs. Instead, to solve the final parsec problem, DM self-interactions would be needed to recreate the spike on a short enough timescale.
@@
Self-interacting dark matter (SIDM) is a class of particle physics models assuming that DM has self-interactions beyond gravity. Initially, SIDM was proposed to solve problems on galactic scales, namely reducing the abundance of Milky Way satellites and the density in the centre of DM halos (Spergel & Steinhardt, 2000). DM self-interactions have been studied in systems covering many orders of magnitude in mass, ranging from dwarf galaxies to galaxy clusters (e.g. Vogelsberger et al., 2014; Fitts et al., 2019; Robertson et al., 2019; Zeng et al., 2022; Ragagnin et al., 2024; Sabarish et al., 2024). Observations of those systems have been used to constrain the strength of the self-interactions (e.g. Sagunski et al., 2021; Despali et al., 2022; Kong et al., 2024; Yang et al., 2024; Zhang et al., 2024). For a review on SIDM, we refer the reader to Tulin & Yu (2018); Adhikari et al. (2022).
@@
When DM is self-interacting, the density profile of the DM spike may change compared to collisionless matter. For SIDM, the slope of the density profile of the DM spike could be smaller than in the case of collisionless DM (Shapiro & Paschalidis, 2014). The lower density would allow relaxing constraints from DM annihilation (Alvarez & Yu, 2021). Importantly, this also affects the dynamical friction, as it depends on the density of the DM spike, and thus also the GW signal (e.g., Macedo et al., 2013; Yue & Han, 2018; Becker et al., 2022). However, given the properties of the density spike, one may still need to calibrate the strength of the dynamical friction with simulations. Kavanagh et al. (2020, 2024) run $N$-body simulations of collisionless DM to measure the Coulomb logarithm for a BH orbiting in a DM spike.
Several studies of dynamical friction for various DM models have been conducted. This includes ultralight DM (Wang & Easther, 2022; Boey et al., 2024), also with self-interactions (Glennon et al., 2024), self-interacting scalar DM (Boudon et al., 2022, 2023a, 2023b), super-fluid DM (Berezhiani et al., 2023), fuzzy DM (Lancaster et al., 2020). In this paper, we investigate the strength of the dynamical friction for a medium consisting of SIDM.
Often the binary BHs have been described in the Newtonian limit, however, depending on their evolution stage this may not be valid. Studies beyond Newtonian gravity have been undertaken by various authors (e.g., Traykova et al., 2023; Montalvo et al., 2024; Mukherjee et al., 2024). For this study, we focus on the Newtonian regime, i.e., when the separation of the two BHs is sufficiently large. This allows us to build upon earlier work and simplifies the numerical description compared to the relativistic case. Extending our study to the relativistic case is left for future work.
The rest of this paper is structured as follows: Section 2 deals with the analytic description of dynamical friction and discusses it in the context of a BH moving through a DM spike. In Section 3, we describe the numerical setup that we use to study dynamical friction in the context of DM self-interactions. This includes the simulation code and the initial conditions and parameters for the simulations. It follows the analysis of our simulations and the results (Section 4). In Section 5, we discuss the limitations of this work as well as directions for further research. Finally, we summarise and conclude in Section 6. Additional information is provided in the appendix.
@@ 2.1 DM spike
Figure 1: The density and one-dimensional velocity dispersion are shown for a DM density spike as a function of radius according to Eq. (2). We assume $r_{\mathrm{sp}}=1\mu\mathrm{pc}$, $\rho_{\mathrm{sp}}=5.45\times 10^{-3}\,\mathrm{M_{\odot}}\,\mu\mathrm{pc}^{-3}$, $\alpha_{\mathrm{sp}}=7/3$, and $M_{\mathrm{BH}}=10^{3}\,\mathrm{M}_{\odot}$. In addition, we indicate the Schwarzschild radius $R_{s}=2\mathrm{G}M/c^{2}$. For computing the velocity dispersion, we follow Kremer (2022). Here, we assume an isotropic velocity distribution. In contrast to Kremer (2022), we solve their eq. 27 including higher-order terms instead of eq. 28. Moreover, we show the results for $m/(k_{\mathrm{B}}T)\approx 6.53\times 10^{3}\,\mathrm{s}^{2}\,\mathrm{km}^{-2}$, with $m$ being the rest mass of a physical DM particle, $k_{\mathrm{B}}$ the Boltzmann constant and $T$ the absolute temperature of the system.
@@
The radius $r_{6}=10^{-6}\,\mathrm{pc}$ is a reference radius at which the density is given by $\rho_{6}$. Note, with this choice, we follow Coogan et al. (2022).
@@
For the relativistic corrections on the properties of the DM spike, we refer the reader to Eda et al. (2015); Tang et al. (2021); Capozziello et al. (2023); John et al. (2023). In Fig. 1 we show the corresponding spike density and velocity dispersion as a function of radius. We note that the density profile of the DM around the BH could be even more complicated when considering a realistic formation scenario of the BH (Bertone et al., 2024).
As we will see in the next subsection the Mach number of the secondary BH moving through the density spike plays a crucial role in determining the dynamical friction force acting on it. The Mach number, $\mathcal{M}=v_{\mathrm{BH}}/c_{\mathrm{s}}$, is given by the velocity, $v_{\mathrm{BH}}$, of the BH and the sound speed, $c_{\mathrm{s}}$. The local speed of sound depends on the adiabatic index $\gamma$, and is given by
@@
To estimate $c_{\mathrm{s}}$, we assume for simplicity that the local velocity distribution obeys a Maxwell-Boltzmann distribution with a one-dimensional velocity dispersion $\sigma_{\mathrm{DM}}$. To estimate the Mach number, $\mathcal{M}$, we assume that the DM spike at sufficiently large radii is well described by Eq. 1.
@@
The adiabatic growth model for a DM spike assuming collisionless DM suggests, for an NFW halo, a spike index of $\alpha_{\mathrm{sp}}=7/3$ (Gondolo & Silk, 1999; Fields et al., 2014; Eda et al., 2015). For SIDM, Shapiro & Paschalidis (2014) derived the DM spike to be $\alpha_{\mathrm{sp}}=7/4$. We note that this is based on the assumption that the self-interaction cross-section has the velocity dependence $\sigma\propto v^{-4}$. Moreover, Shapiro & Paschalidis (2014) assumed that the spike resides in the long-mean-free-path regime, in Section 2.3, we explain that this may often not be the case. This implies that the motion of the inspiralling BH is supersonic when being on a circular orbit (see Eq. 5) given a spike index of $\alpha_{\mathrm{sp}}=7/3$ or $\alpha_{\mathrm{sp}}=7/4$. Even for a smaller value of $\alpha_{\mathrm{sp}}$, it is not feasible to reach the subsonic regime on a circular orbit within the range typically considered for the spike index ($31<\alpha_{\mathrm{sp}}<$ 3). However, a lower value for $\alpha_{\mathrm{sp}}$ might be possible depending on the formation scenario (Ullio et al., 2001).
@@ 2.2 Dynamical friction
The dynamical friction force for a perturber moving through a collisionless medium of constant density was derived by Chandrasekhar (1943). The strongly collisional case with an object moving through a gaseous medium was studied analytically by several authors (e.g., Rephaeli & Salpeter, 1980; Just et al., 1986; Just & Kegel, 1990). Ostriker (1999) derived an analytic description for this regime. Following her, the dynamical friction force can be expressed as
@@
The effective size of the BH is indicated by $r_{\mathrm{min}}$, and $r_{\mathrm{max}}$ denotes the effective size of the surrounding medium. Where, $r_{\mathrm{min}}$, can be understood as the $90\degree$ deflection radius and $r_{\mathrm{max}}$ as the maximum impact parameter (Binney & Tremaine, 2008). We note, that while Chandrasekhar’s formula given by Eq. 7 approximates the dynamical friction well, corrections to the Coulomb logarithm, $\ln(r_{\mathrm{max}}/r_{\mathrm{min}})$, can be necessary (Just & Peñarrubia, 2005; Just et al., 2010), e.g., for velocity distributions deviating from a Maxwell–Boltzmann distribution. For corrections arising from fast moving DM particles on the Chandrasekhar formula, we refer the reader to the work by Dosopoulou (2023) on the dynamical friction in DM spikes.
@@
Note, this equation is based on the assumption that $(c_{s}-v_{\mathrm{BH}})\,t$ exceeds the effective size, $r_{\mathrm{min}}$, of the perturber and that $(c_{s}+v_{\mathrm{BH}})\,t$ is smaller than the effective size, $r_{\mathrm{max}}$, of the surrounding medium. The time, $t$, indicates for how long the BH is perturbing the DM distribution. In the supersonic case with $\mathcal{M}>1$, we have
@@
In Fig. 2, we plot the drag force computed according to Ostriker (1999) for our test setup used in Section 4. We can see that the dynamical friction for velocities slightly above the sound speed is becoming fairly strong in the gaseous case and sharply drops when the velocity falls below the speed of sound.
In the previous subsection, we have seen that the BH would typically move with a supersonic velocity. In consequence, the inspiralling BH will experience enhanced dynamical friction if the DM is not collisionless. However, if the BH is on an eccentric orbit, it might partially, i.e., when it is close to its apocentre, move with a subsonic velocity. We also note, that dynamical friction can contribute to a circularization of the orbit (Becker et al., 2022; Karydas et al., 2024).
In general, the dynamical friction acting on inspiralling BHs may not only arise from the DM spike but also from the accretion disk. The latter can also be described with the Ostriker formula (e.g., Szölgyén et al., 2022; Becker & Sagunski, 2023).
Figure 2: The dynamical friction force is shown as a function of the Mach number, i.e. the velocity in units of the sound speed. The dashed lines indicated the dynamical friction for a collisionless medium according to the Chandrasekhar formula (Chandrasekhar, 1943) and the solid lines are for a gaseous medium as given by Ostriker (1999). Note, for the plot, we assume $r_{\mathrm{max}}=\mathcal{M}\,c_{s}\,t$ and $\ln(r_{\mathrm{max}}/r_{\mathrm{min}})=\ln(\mathcal{M})+C$. The line colours correspond to different values of $C\in\{4.0,6.0,8.0,10.0,12.0\}$. In addition, the red dotted lines indicate the velocity of a secondary BH moving on a circular orbit for different adiabatic indices.

### v2

@@ Abstract
Context. Merging compact objects such as binary black holes provide a promising probe for the physics of dark matter (DM). The gravitational waves emitted during inspiral potentially allow one to detect DM spikes around black holes. This is because the dynamical friction force experienced by the inspiralling black hole alters the orbital period and thus the gravitational wave signal.
@@
Methods. To study the dynamical friction force, we use idealised $N$-body simulations, where we include self-interacting dark matter.
Results. We find that the dynamical friction force for inspiralling black holes would be typically enhanced by DM self-interactions compared to a collisionless medium (ignoring differences in the DM density). At lower velocities below the sound speed, we find that the dynamical friction force can be reduced by the presence of self-interactions.
@@ 1 Introduction
Since the first direct detection of gravitational waves (GWs) from merging stellar-mass black holes (BHs) and neutron stars with the Laser Interferometer Gravitational-Wave Observatory (LIGO; Abbott et al., 2016, 2017a, 2017b, 2019), 100 years after the formulation of Einstein’s theory of general relativity (Einstein, 1916) and about 40 years of the first indirect evidence (Hulse & Taylor, 1975; Taylor et al., 1976; Weisberg et al., 1981), GW astronomy has become an important and very promising field to probe fundamental physics. The GWs emitted by binary mergers do not only allow for testing general relativity at unprecedented levels, but also to probe other physical processes. This in particular includes the physics of neutron stars. Binary mergers of those objects could allow probing the equation of state of ultrahigh-density matter (e.g. Baiotti, 2022). In the future, GW detectors will become sensitive to sub-Hertz GW frequencies, thanks to space-based GW observatories such as the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al., 2017) and Taiji (Hu & Wu, 2017). This will allow us to observe the evolution of intermediate- and extreme-mass ratio inspirals (I/EMRIs), in which a small stellar-mass compact object inspirals into a much larger intermediate/supermassive black hole, over long times and provide a sensitive probe of BH environments.
@@
BHs might be ‘dressed’ with DM, in other words, surrounded by a dense DM spike (e.g. Gondolo & Silk, 1999; Bertone et al., 2005; Bertone, 2024). The first robust evidence for an observation of a DM spike around a BH is claimed to be found by Chan & Lee (2024), see also the work by Alachkar et al. (2023). They studied the supermassive black hole binary system OJ 287 with outburst observations covering a time span of 104 years. The presence of a DM spike can impact the orbit of an inspiralling BH and make the GW signal sensitive to DM physics (e.g. Barausse et al., 2014). This could in principle be detectable, for example for M87* (Daghigh & Kunstatter, 2024). The inspiralling BH is decelerated by the dynamical friction force, which depends on the density and the velocity dispersion of the DM spike (Chandrasekhar, 1943). In consequence, the orbital decay of an inspiralling BH allows probing properties of the DM spike and thus the particle nature of DM.
@@
Potentially, such DM spikes could lead to a faster decay of the orbit at small separations ($\approx 1$ pc). This would provide an explanation for the final parsec problem, that is, how binary BHs can lose enough energy and angular momentum to reach separations small enough that the energy loss due to the emission of GWs is sufficient to make them merge within a sufficiently small timescale (e.g. Milosavljević & Merritt, 2003). In this context, Alonso-Álvarez et al. (2024) claim that collisionless DM is not sufficient to make the BH orbit decay fast enough because the DM spike would be destroyed by the friction energy injected from the BHs. Instead, to solve the final parsec problem, DM self-interactions would be needed to recreate the spike on a short enough timescale.
@@
Self-interacting dark matter (SIDM) is a class of particle physics models assuming that DM has self-interactions beyond gravity. Initially, SIDM was proposed to solve problems on galactic scales, namely reducing the abundance of Milky Way satellites and the density in the centre of DM haloes (Spergel & Steinhardt, 2000). DM self-interactions have been studied in systems covering many orders of magnitude in mass, ranging from dwarf galaxies to galaxy clusters (e.g. Vogelsberger et al., 2014; Fitts et al., 2019; Robertson et al., 2019; Zeng et al., 2022; Ragagnin et al., 2024; Sabarish et al., 2024). Observations of those systems have been used to constrain the strength of the self-interactions (e.g. Sagunski et al., 2021; Despali et al., 2022; Kong et al., 2024; Yang et al., 2024; Zhang et al., 2024). For a review on SIDM, we refer the reader to Tulin & Yu (2018) and Adhikari et al. (2022).
@@
When DM is self-interacting, the density profile of the DM spike may change compared to collisionless matter. For SIDM, the slope of the density profile of the DM spike could be smaller than in the case of collisionless DM (Shapiro & Paschalidis, 2014). The lower density would allow relaxing constraints from DM annihilation (Alvarez & Yu, 2021). Importantly, this also affects the dynamical friction, as it depends on the density of the DM spike, and thus also the GW signal (e.g. Macedo et al., 2013; Yue & Han, 2018; Becker et al., 2022). However, given the properties of the density spike, one may still need to calibrate the strength of the dynamical friction with simulations. Kavanagh et al. (2020, 2024) run $N$-body simulations of collisionless DM to measure the Coulomb logarithm for a BH orbiting in a DM spike.
Several studies of dynamical friction for various DM models have been conducted. This includes ultralight DM (Wang & Easther, 2022; Boey et al., 2024), also with self-interactions (Glennon et al., 2024), self-interacting scalar DM (Boudon et al., 2022, 2023, 2024; Kadota et al., 2024), super-fluid DM (Berezhiani et al., 2024), and fuzzy DM (Lancaster et al., 2020). In this paper, we investigate the strength of the dynamical friction for a medium consisting of SIDM.
Often the binary BHs have been described in the Newtonian limit, however, depending on their evolution stage this may not be valid. Studies beyond Newtonian gravity have been undertaken by various authors (e.g. Traykova et al., 2023; Montalvo et al., 2024; Mukherjee et al., 2024). For this study, we focus on the Newtonian regime, that is, when the separation of the two BHs is sufficiently large. This allows us to build upon earlier work and simplifies the numerical description compared to the relativistic case. Extending our study to the relativistic case is left for future work.
The rest of this paper is structured as follows: Sect. 2 deals with the analytic description of dynamical friction and discusses it in the context of a BH moving through a DM spike. In Sect. 3, we describe the numerical setup that we use to study dynamical friction in the context of DM self-interactions. This includes the simulation code and the initial conditions (ICs) and parameters for the simulations. It follows the analysis of our simulations and the results (Sect. 4). In Sect. 5, we discuss the limitations of this work as well as directions for further research. Finally, we summarise and conclude in Sect. 6. Additional information is provided in the appendix.
@@ 2.1 DM spike
Figure 1: Density and one-dimensional velocity dispersion for a DM density spike as a function of radius according to Eq. (2). We assume $r_{\mathrm{sp}}=1\mu\mathrm{pc}$, $\rho_{\mathrm{sp}}=5.45\times 10^{-3}\,\mathrm{M_{\odot}}\,\mu\mathrm{pc}^{-3}$, $\alpha_{\mathrm{sp}}=7/3$, and $M_{\mathrm{BH}}=10^{3}\,\mathrm{M}_{\odot}$. In addition, we indicate the Schwarzschild radius $R_{s}=2\mathrm{G}M/c^{2}$. For computing the velocity dispersion, we followed Kremer (2022). Here, we assumed an isotropic velocity distribution. In contrast to Kremer (2022), we solved their Eq. 27 including higher-order terms instead of Eq. 28. Moreover, we show the results for $m/(k_{\mathrm{B}}T)\approx 6.53\times 10^{3}\,\mathrm{s}^{2}\,\mathrm{km}^{-2}$, with $m$ being the rest mass of a physical DM particle, $k_{\mathrm{B}}$ the Boltzmann constant and $T$ the absolute temperature of the system.
@@
The radius $r_{6}=10^{-6}\,\mathrm{pc}$ is a reference radius at which the density is given by $\rho_{6}$. We note that with this choice, we follow Coogan et al. (2022).
@@
For the relativistic corrections on the properties of the DM spike, we refer the reader to Eda et al. (2015), Tang et al. (2021), Capozziello et al. (2023), and John et al. (2024). In Fig. 1 we show the corresponding spike density and velocity dispersion as a function of radius. We note that the density profile of the DM around the BH could be even more complicated when considering a realistic formation scenario of the BH (Bertone et al., 2024).
As we see in the next subsection the Mach number of the secondary BH moving through the density spike plays a crucial role in determining the dynamical friction force acting on it. The Mach number, $\mathcal{M}=v_{\mathrm{BH}}/c_{\mathrm{s}}$, is given by the velocity, $v_{\mathrm{BH}}$, of the BH and the sound speed, $c_{\mathrm{s}}$. The local speed of sound depends on the adiabatic index $\gamma$, and is given by
@@
To estimate $c_{\mathrm{s}}$, we assume for simplicity that the local velocity distribution obeys a Maxwell-Boltzmann distribution with a one-dimensional velocity dispersion $\sigma_{\mathrm{DM}}$. For the Mach number, $\mathcal{M}$, we make the assumption that the DM spike at sufficiently large radii is well described by Eq. 1.
@@
The adiabatic growth model for a DM spike assuming collisionless DM suggests, for an NFW halo, a spike index of $\alpha_{\mathrm{sp}}=7/3$ (Gondolo & Silk, 1999; Fields et al., 2014; Eda et al., 2015). For SIDM, Shapiro & Paschalidis (2014) derived the DM spike to be $\alpha_{\mathrm{sp}}=7/4$. We note that this is based on the assumption that the self-interaction cross-section has the velocity dependence $\sigma\propto v^{-4}$. Moreover, Shapiro & Paschalidis (2014) assumed that the spike resides in the long-mean-free-path regime, in Sect. 2.3, we explain that this may eventually not always be the case. This implies that the motion of the inspiralling BH is supersonic when being on a circular orbit (see Eq. 5) given a spike index of $\alpha_{\mathrm{sp}}=7/3$ or $\alpha_{\mathrm{sp}}=7/4$. Even for a smaller value of $\alpha_{\mathrm{sp}}$, it is not feasible to reach the subsonic regime on a circular orbit within the range typically considered for the spike index ($31<\alpha_{\mathrm{sp}}<$ 3). However, a lower value for $\alpha_{\mathrm{sp}}$ might be possible depending on the formation scenario (Ullio et al., 2001).
@@ 2.2 Dynamical friction
The dynamical friction force for a perturber moving through a collisionless medium of constant density was derived by Chandrasekhar (1943). The strongly collisional case with an object moving through a gaseous medium was studied analytically by several authors (e.g. Rephaeli & Salpeter, 1980; Just et al., 1986; Just & Kegel, 1990). Ostriker (1999) derived an analytic description for this regime. Following her, the dynamical friction force can be expressed as
@@
The effective size of the BH is indicated by $r_{\mathrm{min}}$, and $r_{\mathrm{max}}$ denotes the effective size of the surrounding medium. Where, $r_{\mathrm{min}}$, can be understood as the $90\degree$ deflection radius and $r_{\mathrm{max}}$ as the maximum impact parameter (Binney & Tremaine, 2008). We note that while Chandrasekhar’s formula given by Eq. 7 approximates the dynamical friction well, corrections to the Coulomb logarithm, $\ln(r_{\mathrm{max}}/r_{\mathrm{min}})$, can be necessary (Just & Peñarrubia, 2005; Just et al., 2010), for example for velocity distributions deviating from a Maxwell–Boltzmann distribution. For corrections arising from fast moving DM particles on the Chandrasekhar formula, we refer the reader to the work by Dosopoulou (2023) on the dynamical friction in DM spikes.
@@
We note that this equation is based on the assumption that $(c_{s}-v_{\mathrm{BH}})\,t$ exceeds the effective size, $r_{\mathrm{min}}$, of the perturber and that $(c_{s}+v_{\mathrm{BH}})\,t$ is smaller than the effective size, $r_{\mathrm{max}}$, of the surrounding medium. The time, $t$, indicates for how long the BH is perturbing the DM distribution. In the supersonic case with $\mathcal{M}>1$, we have
@@
In Fig. 2, we plot the drag force computed according to Ostriker (1999) for our test setup used in Sect. 4. We can see that the dynamical friction for velocities slightly above the sound speed is becoming fairly strong in the gaseous case and sharply drops when the velocity falls below the speed of sound.
In the previous subsection, we have seen that the BH would typically move with a supersonic velocity. In consequence, the inspiralling BH experiences enhanced dynamical friction if the DM is not collisionless. However, if the BH is on an eccentric orbit, it might partially, when it is close to its apocentre, move with a subsonic velocity. We also note that dynamical friction can contribute to a circularisation of the orbit (Becker et al., 2022; Karydas et al., 2024).
In general, the dynamical friction acting on inspiralling BHs may not only arise from the DM spike but also from the accretion disk. The latter can also be described with the Ostriker formula (e.g. Szölgyén et al., 2022; Becker & Sagunski, 2023).
Figure 2: Dynamical friction force as a function of the Mach number, i.e. the velocity in units of the sound speed. The dashed lines indicated the dynamical friction for a collisionless medium according to the Chandrasekhar formula (Chandrasekhar, 1943) and the solid lines are for a gaseous medium as given by Ostriker (1999). We note that for the plot, we assume $r_{\mathrm{max}}=\mathcal{M}\,c_{s}\,t$ and $\ln(r_{\mathrm{max}}/r_{\mathrm{min}})=\ln(\mathcal{M})+C$. The line colours correspond to different values of $C\in\{4.0,6.0,8.0,10.0,12.0\}$. In addition, the red dotted lines indicate the velocity of a secondary BH moving on a circular orbit for different adiabatic indices.